OpenStudy (loginusernamed):

Solve for N. (4 2/3 + 2 4/7) + 7 3/7 = N+12

OpenStudy (anonymous):

Use\[BODMAS~rule\]where\[B=bracket\]\[O=of\]\[D=division\]\[M=multiplication\]\[A=addition\]\[S=subtraction\]

OpenStudy (anonymous):

\[\color{red}{(4\frac{ 2 }{ 3 }+2\frac{ 4 }{ 7 })}+7\frac{ 3 }{ 7 }=N+12\]Make the denominator same

OpenStudy (anonymous):

\[\color{red}{(4\frac{ 2\times7 }{ 3\times7 }+2\frac{ 4\times3 }{ 7\times3 })}+7\frac{ 3 }{ 7 }=N+12\]